# CliCARE: Grounding Large Language Models in Clinical Guidelines for Decision Support over Longitudinal Cancer Electronic Health Records

### v1

@@ Abstract
Large Language Models (LLMs) hold significant promise for improving clinical decision support and reducing physician burnout by synthesizing complex, longitudinal cancer Electronic Health Records (EHRs). However, their implementation in this critical field faces three primary challenges: the inability to effectively process the extensive length and multilingual nature of patient records for accurate temporal analysis; a heightened risk of clinical hallucination, as conventional grounding techniques such as Retrieval-Augmented Generation (RAG) do not adequately incorporate process-oriented clinical guidelines; and unreliable evaluation metrics that hinder the validation of AI systems in oncology. To address these issues, we propose CliCARE, a framework for Grounding Large Language Models in Clinical Guidelines for Decision Support over Longitudinal CAncer Electronic Health REcords. The framework operates by transforming unstructured, longitudinal EHRs into patient-specific Temporal Knowledge Graphs (TKGs) to capture long-range dependencies, and then grounding the decision support process by aligning these real-world patient trajectories with a normative guideline knowledge graph. This approach provides oncologists with evidence-grounded decision support by generating a high-fidelity clinical summary and an actionable recommendation. We validated our framework using large-scale, longitudinal data from a private Chinese cancer dataset and the public English MIMIC-IV dataset. In these diverse settings, CliCARE significantly outperforms strong baselines, including leading long-context LLMs and Knowledge Graph-enhanced RAG methods. The clinical validity of our results is supported by a robust evaluation protocol, which demonstrates a high correlation with assessments made by expert oncologists.
@@ 1 Introduction
Large Language Models (LLMs) are emerging as promising tools for clinical decision support, with current research demonstrating their potential to serve as collaborative partners that augment expert workflows, alleviate clinician workloads, and improve decision-making in complex fields such as oncology (Hager et al. 2024; Rajashekar et al. 2024). However, their integration into high-stakes clinical practice is far from straightforward. The reality of clinical oncology involves physicians navigating immense cognitive burdens from manually integrating fragmented data within multi-year Electronic Health Records (EHRs), a key contributor to professional burnout (Warner et al. 2020; Sinsky et al. 2016). This chaotic, unstructured environment amplifies the critical disparity between LLMs’ high performance on standardized benchmarks and their actual capabilities in the clinic. Indeed, systematic reviews indicate that their performance in cancer decision-making is inconsistent with critical safety aspects frequently unaddressed (Hao et al. 2025), while other recent studies confirm that even state-of-the-art models struggle to adhere to treatment guidelines or accurately interpret laboratory results (Hager et al. 2024). Therefore, the frontier of this field is not merely the development of more powerful models, but the creation of robust frameworks that ensure these technologies are reliable, safe, and effectively grounded in expert medical knowledge to truly augment, not supplant, the role of the physician.In practice, augmenting the expert physician’s role means supporting their core clinical workflow: synthesizing a patient’s multi-year history into a coherent Clinical Summary, and from that summary, generating an actionable Clinical Recommendation for future treatment.
@@
However, automating this expert workflow with existing LLMs faces three fundamental challenges. First, LLMs exhibit a significant inability to perform effective temporal reasoning over the extensive data records typical of cancer EHRs. Our research addresses a corpus containing large samples of patient records, where a single patient’s history can span years, exceed 20,000 tokens, and even include multilingual entries, making brute-force approaches inefficient (Liu et al. 2024). The second challenge is the unacceptable risk of clinical hallucination, which undermines the potential for reliable decision support. Factually incorrect recommendations pose a direct threat to patient safety, a risk that is exacerbated by the limitations of standard Retrieval-Augmented Generation methods. The retrieval of fragmented text fails to capture the sequential dependencies in a patient’s trajectory and cannot effectively bridge the gap with process-oriented clinical guidelines (Li et al. 2024). Finally, the field confronts two interconnected barriers to real-world adoption. The deployment dilemma centers on a fundamental trade-off. On one hand, powerful, closed-source models offer state-of-the-art performance but raise significant concerns regarding cost and patient data privacy. On the other hand, open-source alternatives are more efficient and easier to deploy locally, though they often lag in capability. This trade-off is compounded by the significant challenge of reliable evaluation, as the high-stakes nature of clinical content renders conventional automated metrics untrustworthy, thereby hindering reliable progress and diminishing clinical trust (Wang et al. 2023; Zheng et al. 2023).
@@
To address these barriers, we propose CliCARE, a framework for Grounding Large Language Models in Clinical Guidelines for Decision Support over Longitudinal CAncer Electronic Health REcords. CliCARE first tackles long-context temporal analysis by structuring raw EHRs into Temporal Knowledge Graphs (TKGs) to make temporal relationships explicit (Sec 3.1). It then mitigates hallucinations by grounding the model through a deep alignment of patient trajectories with clinical guidelines (Sec 3.2). This structured representation provides both gold-standard data for fine-tuning specialist models and rich context for large generalist models. Finally, we ensure reliable evaluation via a human-validated LLM-as-a-Judge protocol whose ratings highly correlate with expert judgments (Sec 4.1). Our source code is provided in the supplementary material for reproducibility.
@@
  • We introduce CliCARE, an end-to-end framework that grounds LLMs by transforming EHRs into TKGs and aligning them with clinical guidelines, featuring an adaptable architecture for both generalist and specialist models.
  • We propose a reliable evaluation methodology using a Human-Validated LLM-as-a-Judge, whose ratings show strong Spearman’s correlation with expert oncologists, addressing the limitations of standard automated metrics.
  • Extensive experiments on diverse datasets show CliCARE significantly outperforms robust baselines, while ablation studies confirm the contribution of each component.
Figure 2: A comparison of RAG approaches for long-form longitudinal clinical tasks. (a) Standard RAG often suffers from missing key information and hallucinations. (b) KG-enhanced RAG struggles to model temporal dependencies in patient journeys. (c) In contrast, our CliCARE framework transforms EHRs into Temporal Knowledge Graphs, aligns patient trajectories with guidelines, and generates answers using a distilled specialist model, which are then assessed by our proposed evaluation approach.
@@ 2 Related Work
2.1 LLMs with Long-Form EHRs
Applying LLMs to long-form EHRs for clinical decision support is fundamentally constrained by the challenge of long-context processing. The evaluation of these challenges has been systematized through benchmarks such as LongBench (Bai et al. 2024), with specialized medical benchmarks like MedOdyssey emerging to assess these capabilities in a clinical context (Fan et al. 2024). Prominent issues include the ”lost-in-the-middle” problem and the degradation of performance during domain-specific fine-tuning (Liu et al. 2024; Zhang et al. 2024; Yang et al. 2024). While economical solutions like Parameter-Efficient Fine-Tuning (PEFT) show promise, these technical advances alone are often insufficient for achieving clinically meaningful outcomes without structured knowledge to guide the model (Dong et al. 2024; Zhang et al. 2023; Nazary et al. 2024). This need has led to a primary strategy of transforming unstructured data into structured formats, a critical step for the robust temporal analysis necessary for summarizing patient journeys. However, while methods such as Patient Journey Knowledge Graphs (PJKGs) exist (Khatib et al. 2025), their variable accuracy and efficiency present reliability challenges for downstream decision support tasks.
@@ 2.2 Knowledge Graph-enhanced LLMs and RAG
Augmenting LLMs with external Knowledge Graphs (KGs) is a crucial strategy for mitigating factual errors and hallucinations, which is essential for safety in high-stakes domains such as healthcare (Khan, Wu, and Chen 2024). This approach helps bridge the gap between general-purpose models and specialized clinical knowledge (Yu and McQuade 2025). However, standard Retrieval-Augmented Generation (RAG) retrieves isolated text snippets, overlooking the complex relational structures necessary for effective clinical decision-making (Lewis et al. 2020; Liu et al. 2024). This limitation has prompted the development of Graph-Aware RAG, which retrieves structured subgraphs instead of disconnected text. Frameworks such as MedRAG, GNN-RAG, and KG2RAG utilize this method to enhance model performance with domain-specific KGs (Zhao et al. 2025; Feng et al. 2024; e Shi et al. 2024). A more advanced frontier moves beyond retrieval to the alignment and fusion of KGs and LLMs at the representation level, projecting both into a unified semantic space for more nuanced, topology-aware outputs (Jiang et al. 2024). This synergy creates a virtuous cycle, where KGs not only ground LLMs, but LLMs are also increasingly used to construct and enrich KGs from text (Maushagen et al. 2024).
@@ 2.3 Assessment of Open-Ended Clinical Generation Tasks
Evaluating open-ended generation from LLMs in high-stakes medical domains presents a critical challenge. Traditional automated metrics, such as ROUGE and BLEU, are widely regarded as inadequate because their emphasis on lexical overlap fails to capture essential aspects like clinical validity, factual accuracy, and safety (Wang et al. 2023; Singhal et al. 2023). In response, research has increasingly focused on more nuanced evaluation methods, including dynamic agent-based assessments (Tu et al. 2024; Tadevosyan et al. 2025) and the scalable ”LLM-as-a-Judge” paradigm (Zheng et al. 2023). However, the reliability of LLM judges is compromised by known systematic biases like positional bias and verbosity, raising significant safety concerns in a field where deep domain expertise is essential (Zheng et al. 2023; Wang et al. 2024). This highlights the urgent need for rigorous methodologies to validate automated judgments against human expert reasoning.
Existing research has treated long-context processing, knowledge grounding, and reliable evaluation as distinct challenges. A significant research gap exists in developing a solution that simultaneously addresses the long-context limitations in real-world EHRs, provides deep grounding in clinical guidelines that exceeds standard RAG, and guarantees trustworthy assessment. CliCARE is designed to bridge this gap by integrating these capabilities into a unified, cohesive pipeline.
@@ 3 CliCARE
In this section, we present the CliCARE framework, as illustrated in Figure 2. CliCARE is designed to systematically analyze long-form, unstructured cancer EHRs to generate a clinically grounded clinical summary and clinical recommendations. A key feature of this design is its extensibility; the guideline knowledge graph can be efficiently updated and expanded to accommodate new clinical evidence and corresponding guidelines.
@@ 3.1 EHR-to-TKGs Transformation
Here, $E_{p}$ represents a structured sequence of key clinical events. Specifically, the core of $f_{\text{pipeline}}$ is an extractive summarization module based on the Longformer model. Given the computational cost of processing the entire $D_{p}$, we partition the document sequence into the most recent clinical note, $d_{\tau_{n}}$, and the historical records, $D_{p}^{\text{hist}}=(d_{\tau_{1}},\dots,d_{\tau_{n-1}})$. We utilize a Longformer model, $\mathcal{M}_{\text{LF}}$, pre-trained on clinical text, to process the extensive historical records $D_{p}^{\text{hist}}$:
@@
where $\theta_{\text{LF}}$ are the model parameters and $S_{p}^{\text{hist}}$ is a summary text that includes the most informative sentences from the historical records. This summary effectively functions as the patient’s ”past medical history.” The most recent clinical note, $d_{\tau_{n}}$, is regarded as the ”history of present illness.” The final structured event sequence $E_{p}$ is created by concatenating these two components, thus providing a chronologically coherent and condensed patient history. From this combined text, key clinical facts—such as diagnostic confirmations, staging updates, treatment regimens, biomarker trends, and imaging assessments—are identified and organized into discrete events (Yang, Wang, and Li 2021; Huang, Altosaar, and Ranganath 2019).
@@
Extracted event sequences $E_{p}$ are organized into a patient-centric Clinical TKG, denoted as $G_{t}=(E_{t},R_{t},T)$, where $E_{t}$ is the set of entities, $R_{t}$ the set of relations, and $T$ the set of timestamps. To enrich the TKG with standardized medical knowledge, we first construct a general, static biomedical knowledge graph $G_{B}=(\mathcal{E}_{B},\mathcal{R}_{B})$, where $\mathcal{E}_{B}$ contains standardized medical concepts and $\mathcal{R}_{B}$ represents the relations between them. For each patient, letting $\mathcal{E}_{p}$ be the set of raw clinical entities extracted from the patient’s record, we instantiate a personalized TKG $G_{t}$ by linking extracted clinical entities from the patient’s record to the concepts in $G_{B}$. This is achieved via an entity linking function $\phi:\mathcal{E}_{p}\rightarrow\mathcal{E}_{B}$, which maps textual mentions in the EHR to their corresponding canonical entries in the biomedical ontology. Each entity $e\in\mathcal{E}_{t}$ in the TKG is a spatiotemporal instance represented as $e=(e_{B},\tau,A)$, where $e_{B}\in\mathcal{E}_{B}$ is the linked standard entity, $\tau\in\mathcal{T}$ is the event timestamp, and $A$ is a set of event-specific attributes.
The TKG employs a hierarchical timestamp granularity by assigning precise timestamps $\tau\in T$ only to macro-level clinical encounters, while linking intra-encounter events through relative temporal relations, thereby mirroring the structure of real-world clinical records to designed to capture capture the dynamic evolution of a patient’s disease course.
Figure 3: Our Trajectory-Guideline Alignment workflow. It fuses patient data with guidelines via semantic matching, LLM-based re-ranking, and iterative bootstrapping expansion to create a comprehensive, evidence-grounded mapping.

### v2

@@ Abstract
Large Language Models (LLMs) hold significant promise for improving clinical decision support and reducing physician burnout by synthesizing complex, longitudinal cancer Electronic Health Records (EHRs). However, their implementation in this critical field faces three primary challenges: the inability to effectively process the extensive length and fragmented nature of patient records for accurate temporal analysis; a heightened risk of clinical hallucination, as conventional grounding techniques such as Retrieval-Augmented Generation (RAG) do not adequately incorporate process-oriented clinical guidelines; and unreliable evaluation metrics that hinder the validation of AI systems in oncology. To address these issues, we propose CliCARE, a framework for Grounding Large Language Models in Clinical Guidelines for Decision Support over Longitudinal CAncer Electronic Health REcords. The framework operates by transforming unstructured, longitudinal EHRs into patient-specific Temporal Knowledge Graphs (TKGs) to capture long-range dependencies, and then grounding the decision support process by aligning these real-world patient trajectories with a normative guideline knowledge graph. This approach provides oncologists with evidence-grounded decision support by generating a high-fidelity clinical summary and an actionable recommendation. We validated our framework using large-scale, longitudinal data from a private Chinese cancer dataset and the public English MIMIC-IV dataset. In these settings, CliCARE significantly outperforms baselines, including leading long-context LLMs and Knowledge Graph-enhanced RAG methods. The clinical validity of our results is supported by a robust evaluation protocol, which demonstrates a high correlation with assessments made by oncologists.
@@ 1 Introduction
Large Language Models (LLMs) are emerging as promising tools for clinical decision support, with current research demonstrating their potential to serve as collaborative partners that augment expert workflows, alleviate clinician workloads, and improve decision-making in complex fields such as oncology (Hager et al. 2024; Rajashekar et al. 2024). However, their integration into high-stakes clinical practice is far from straightforward. The reality of clinical oncology involves physicians navigating cognitive burdens from manually integrating fragmented data within multi-year Electronic Health Records (EHRs), a key contributor to professional burnout (J. L. Warner, D. Dymshyts, G. Hripcsak, and R. S. Miller (2020); 2). This chaotic, unstructured environment amplifies the critical disparity between LLMs’ high performance on standardized benchmarks and their capabilities in the clinic. Indeed, systematic reviews indicate that their performance in cancer decision-making is inconsistent, with critical safety aspects frequently unaddressed (Hao et al. 2025), while other recent studies confirm that even state-of-the-art models struggle to adhere to treatment guidelines or accurately interpret laboratory results (Hager et al. 2024). Therefore, the frontier of this field is not merely the development of more powerful models, but the creation of robust frameworks that ensure these technologies are reliable, safe, and effectively grounded in expert medical knowledge to augment, not supplant, the role of the physician. In practice, augmenting the expert physician’s role means supporting their core clinical workflow: synthesizing a patient’s multi-year history into a coherent Clinical Summary, and from that summary, generating an actionable Clinical Recommendation for future treatment.
@@
However, automating this expert workflow with existing LLMs faces three challenges. First, LLMs exhibit an inability to perform effective temporal reasoning over the extensive data records typical of cancer EHRs. Our research addresses a corpus containing large samples of patient records, where a single patient’s history can span years, exceed 20,000 tokens, and even include multilingual entries, making brute-force approaches inefficient (Liu et al. 2024). The second challenge is the unacceptable risk of clinical hallucination, which undermines the potential for reliable decision support. Factually incorrect recommendations pose a threat to patient safety, a risk that is exacerbated by the limitations of standard Retrieval-Augmented Generation methods. The retrieval of fragmented text fails to capture the sequential dependencies in a patient’s trajectory and cannot effectively bridge the gap with process-oriented clinical guidelines (Li et al. 2024). Finally, the field confronts two interconnected barriers to real-world adoption. The deployment dilemma centers on a trade-off. On one hand, powerful, closed-source models offer state-of-the-art performance but raise significant concerns regarding cost and patient data privacy. On the other hand, open-source alternatives are more efficient and easier to deploy locally, though they often lag in capability. This trade-off is compounded by the significant challenge of reliable evaluation, as the high-stakes nature of clinical content renders conventional automated metrics untrustworthy, thereby hindering reliable progress and diminishing clinical trust (Wang et al. 2023; Zheng et al. 2023).
@@
To address these barriers, we propose CliCARE, a framework for Grounding Large Language Models in Clinical Guidelines for Decision Support over Longitudinal CAncer Electronic Health REcords. CliCARE first tackles long-context temporal analysis by structuring raw EHRs into Temporal Knowledge Graphs (TKGs) to make temporal relationships explicit (Sec 3.1). It then mitigates hallucinations by grounding the model through a deep alignment of patient trajectories with clinical guidelines (Sec 3.2). This representation provides both guideline-grounded data for fine-tuning specialist models and rich context for large generalist models. Finally, we ensure reliable evaluation via an Expert-Validated LLM-as-a-Judge protocol whose ratings correlate with expert judgments (Sec 4.1).
@@
  • We introduce CliCARE, an end-to-end framework that grounds LLMs by transforming unstructured clinical text from EHRs into TKGs and aligning them with clinical guidelines. It features an adaptable architecture for both generalist and specialist models.
  • We propose a reliable evaluation methodology using an Expert-Validated LLM-as-a-Judge, whose ratings are highly correlated with expert oncologists, addressing the limitations of automated metrics.
  • Extensive experiments on diverse datasets show CliCARE significantly outperforms baselines, while ablation studies confirm the contribution of each component.
Figure 2: A comparison of RAG approaches for long-form longitudinal clinical tasks. (a) Standard RAG suffers from missing key information and hallucinations. (b) KG-enhanced RAG struggles to model temporal dependencies in patient journeys. (c) In contrast, our CliCARE framework transforms EHRs into Temporal Knowledge Graphs, aligns patient trajectories with guidelines, and generates answers using a distilled specialist model, which are then assessed by our evaluation approach.
@@ 2 Related Work
2.1 LLMs for Decision Support with Long-Form EHRs
Leveraging LLMs for clinical decision support holds promise for synthesizing complex EHRs. However, the longitudinal, long-form nature of these records presents a barrier. Models often struggle with long-context processing, leading to issues like the lost-in-the-middle problem and performance degradation (Liu et al. 2024; Zhang et al. 2025; Yang et al. 2025b). These challenges are now being systematically evaluated by specialized benchmarks like LongBench and MedOdyssey (Bai et al. 2024; Fan et al. 2024).
To mitigate these issues and unlock the potential of LLMs for decision support, a primary strategy is to transform unstructured data into structured formats. Current approaches range from prompting LLMs on pre-structured data for prediction tasks (Zhu et al. 2024b), or enhancing predictions from structured codes via retrieval augmentation (Xu et al. 2024), to generating Patient Journey Knowledge Graphs (PJKGs) from raw text, though the latter can face reliability challenges (Khatib et al. 2025). The emergence of sophisticated, multi-agent architectures like ColaCare further raises the bar for comprehensive EHR modeling (Wang et al. 2025), highlighting the field’s push towards robust, structured solutions for decision support.
@@ 2.2 Knowledge Graph-enhanced LLMs and RAG
Augmenting LLMs with external Knowledge Graphs (KGs) is a crucial strategy for mitigating factual errors and hallucinations, which is essential for safety in high-stakes domains such as healthcare (Khan et al. 2024). This approach helps bridge the gap between general-purpose models and specialized clinical knowledge (Yu and McQuade 2025). However, standard RAG often retrieves isolated text snippets, overlooking the relational structures necessary for clinical decision-making (Lewis et al. 2020). Recent efforts seek to address this by focusing retrieval on specific knowledge sources, such as directly incorporating clinical practice guidelines (Oniani et al. 2024) or enhancing multimodal EHR analysis through RAG-driven frameworks like REALM (Zhu et al. 2024a). To more fundamentally resolve knowledge fragmentation, Graph-Aware RAG moves beyond text retrieval to extract structured information, retrieving entire knowledge paths to improve diagnosis prediction (Gao et al. 2025) or pulling relevant subgraphs via frameworks like MedRAG and GNN-RAG (Zhao et al. 2025; Mavromatis and Karypis 2025). An even more advanced frontier moves beyond retrieval to the alignment and fusion of KGs and LLMs at the representation level (Jiang et al. 2024), a synergy that creates a virtuous cycle of grounding and enrichment (Maushagen et al. 2024).
@@ 2.3 Assessment of Open-Ended Clinical Generation Tasks
Evaluating open-ended generation from LLMs in high-stakes medical domains presents a challenge. Traditional automated metrics, such as ROUGE and BLEU, are regarded as inadequate because their emphasis on lexical overlap fails to capture essential aspects like clinical validity, factual accuracy, and safety (Wang et al. 2023; Singhal et al. 2023). In response, research has increasingly focused on more nuanced evaluation methods, including dynamic agent-based assessments (Tu et al. 2024; Tadevosyan et al. 2025) and the scalable LLM-as-a-Judge paradigm (Zheng et al. 2023). However, the reliability of LLM judges is compromised by known systematic biases like positional bias and verbosity, raising safety concerns in a field where deep domain expertise is essential (Zheng et al. 2023; Wang et al. 2024). This highlights the urgent need for rigorous methodologies to validate automated judgments against human expert reasoning.
Existing research has treated long-context processing, knowledge grounding, and reliable evaluation as distinct challenges. A research gap exists in developing a solution that simultaneously addresses the long-context limitations in real-world EHRs, provides deep grounding in clinical guidelines that exceeds standard RAG, and guarantees trustworthy assessment. CliCARE is designed to bridge this gap by integrating these capabilities into a unified pipeline.
@@ 3 CliCARE
In this section, we present the CliCARE framework, as illustrated in Figure 2. CliCARE is designed to systematically analyze long-form, unstructured cancer EHRs to generate a clinically grounded clinical summary and clinical recommendations. A key feature of this design is its extensibility: the guideline knowledge graph can be efficiently updated and expanded to accommodate new clinical evidence and corresponding guidelines.
@@ 3.1 EHR-to-TKGs Transformation
Here, $E_{p}$ represents a structured sequence of key clinical events. Specifically, the core of $f_{\text{pipeline}}$ is an extractive summarization module based on the Longformer model. Given the computational cost of processing the entire $D_{p}$, we partition the document sequence into the most recent clinical note, $d_{\tau_{n}}$, and the historical records, $D_{p}^{\text{hist}}=(d_{\tau_{1}},\dots,d_{\tau_{n-1}})$. We utilize a Longformer model (Beltagy et al. 2020), $\mathcal{M}_{\text{LF}}$, pre-trained on clinical text, to process the extensive historical records $D_{p}^{\text{hist}}$:
@@
where $\theta_{\text{LF}}$ are the model parameters and $S_{p}^{\text{hist}}$ is a summary text that includes the most informative sentences from the historical records. This summary effectively functions as the patient’s past medical history. The most recent clinical note, $d_{\tau_{n}}$, is regarded as the history of the present illness. The final structured event sequence $E_{p}$ is created by concatenating these two components, thus providing a chronologically coherent and condensed patient history. From this combined text, key clinical facts—such as diagnostic confirmations, staging updates, treatment regimens, biomarker trends, and imaging assessments—are identified and organized into discrete events using BERT for information extraction(Yang et al. 2021; Huang et al. 2019).
@@
Extracted event sequences $E_{p}$ are organized into a patient-centric Clinical TKG, denoted as $G_{t}=(E_{t},R_{t},T)$, where $E_{t}$ is the set of entities, $R_{t}$ the set of relations, and $T$ the set of timestamps. To enrich the TKG with standardized medical knowledge, we first construct a general, static biomedical knowledge graph $G_{B}=(\mathcal{E}_{B},\mathcal{R}_{B})$, where $\mathcal{E}_{B}$ contains standardized medical concepts and $\mathcal{R}_{B}$ represents the relations. For each patient, letting $\mathcal{E}_{p}$ be the set of raw clinical entities extracted from the patient’s record, we instantiate a personalized TKG $G_{t}$ by linking extracted clinical entities from the patient’s record to the concepts in $G_{B}$. This uses an entity linking function $\phi:\mathcal{E}_{p}\rightarrow\mathcal{E}_{B}$, which maps textual mentions in the EHR to their corresponding canonical entries in the biomedical ontology. Each entity $e\in\mathcal{E}_{t}$ is represented as $e=(e_{B},\tau,A)$, where $e_{B}\in\mathcal{E}_{B}$ is the linked standard entity, $\tau\in T$ is the event timestamp, and $A$ is a set of event-specific attributes.
The TKG employs a hierarchical timestamp granularity by assigning precise timestamps $\tau\in T$ only to macro-level clinical encounters, while linking intra-encounter events through relative temporal relations, thereby mirroring the structure of real-world clinical records designed to capture the dynamic evolution of a patient’s disease course.
Figure 3: Trajectory-Guideline Alignment workflow. It fuses patient data with guidelines via semantic matching, LLM-based Reranking, and iterative bootstrapping expansion to create a comprehensive, evidence-grounded mapping.